OpenStudy (anonymous):

PLEASE HELP! Solve each equation for x. assume that a is not equal to 0. ax+7=5

OpenStudy (anonymous):

ax+7=5

OpenStudy (anonymous):

\[\frac{ 1 }{ a }x-4=9\]

OpenStudy (anonymous):

2-ax=-8

OpenStudy (anonymous):

I'm having troubles with all three of these problems.. please help me!! I don't understand them!